A machine tool cylinder is connected regeneratively to give a rapid approach speed of 10 m/min for a stroke of 1 m with a theoretical thrust of 2.5 tonnes. It is then switched to conventional connection to provide a pressing speed of 0.25 m/min for 0.5 m with a theoretical thrust of 10 tonnes. The maximum pressure at the cylinder is to be 200 bar.


(a) Select a suitable standard size metric cylinder.


(b) Calculate the pump delivery required for both parts of the extend stroke.


1
Expert's answer
2021-12-21T00:33:01-0500

Total thrust "F= 2.5 T = 24516 N"

maximum pressure "p = 200 bar = 2 0 \frac{N}{m^2}"

"A = \frac{\pi}{4} d^2 = \frac{24516}{20}=d=39.50" mm

Piston velocity during extension stroke

"P_E = \frac{Q}{A_p}=\frac{200}{\frac{\pi}{4}39.50^2}=0.1632 \frac{m}{s}"